Use proof by contraposition to show that if x + y ≥ 2, where x and y are real numbers, then x ≥ 1 or y ≥ 1.
1
Expert's answer
2020-10-26T18:57:44-0400

"\text{Let }x<1 \text{and }y<1"

"\text{the sum of inequalities is also a true inequality that is}"

"x+y<2"

"\text{If } (x<1 )\land(y<1)\text{ then}\sum{(x+y)}<2"

"\text{The contrapositive version:}"

"\text{if }\sum{(x+y)}\ge2;\text{ then}¬((x<1)\land(y<1))"

"¬(x<1) \text{ is }(x\ge1);\ ¬(y<1) \text{ is }(y\ge1);"

"\text{according to de Morgan’s laws:}"

"¬((x<1)\land(y<1))=¬(x<1)\lor¬(y<1)=(x\ge1)\lor(y\ge1)"

"\text{if }\sum{(x+y)}\eqslantgtr2\text{ then}(x\ge)\text{ or }(y\ge1)"

"\text{proof done}"